OpenStudy (anonymous):

determine whether the function is periodic?if it is periodic,find the smallest (fundamental) period. a) sin x - cos [(2)^1/2]x

OpenStudy (alfie):

This function is not periodic, so we have no period.

OpenStudy (anonymous):

how u know that its not periodic?teach me please..

OpenStudy (anonymous):

\[\sin x - \cos \sqrt{2}x\]

OpenStudy (alfie):

http://tutorial.math.lamar.edu/Classes/DE/PeriodicOrthogonal.aspx Try to read this, is quite easy to understand.

OpenStudy (anonymous):

ouh..then how about this? \[\cos 3x-\sin 7x\]

OpenStudy (alfie):

Both sine and cosine are periodic in your last function, a sum of two periodic functions is a periodic function

OpenStudy (anonymous):

hmm..can u explain in detail..im sorry if im asking too much..

OpenStudy (alfie):

\[\cos3x - sen7x = \cos(3x+T) - \sin(7x + T)\] We know that both cosine and sine period happens to be \[2\pi\] You try to plug in \[2\pi\] inside of the given function and you'll see you'll have the very same value, since it makes a 360° spin and comes back to the starting point.

OpenStudy (anonymous):

meaning that when we substitute T to \[2\pi\] it will give the same answer..is that what u mean?

OpenStudy (alfie):

Yes, as a simple example. \[\sin(90°)\] is eqaul to 1. Now, if you do: \[\sin(90°) = \sin(90° + T) \rightarrow \sin(90°) = \sin(90° + 360°)\] Obviously the period of the sin is 360°, so is sin of 360°+90 = sin90° ? Try it yourself, how much is \[\sin(450°)\] ? If it is 1 as well, we have a periodic function.

OpenStudy (anonymous):

ouhh..now i understand..thank u so much for helping me !! ^^,

OpenStudy (alfie):

You're welcome :)

OpenStudy (anonymous):

^^,

OpenStudy (anonymous):

alfie how come you say sin x - cos (sqrt x ) is not periodic?

OpenStudy (alfie):

Take a look at the graph.